# Efficient Monte Carlo Event Generation for Neutrino-Nucleus Exclusive Cross Sections

### v2

@@ I Introduction
To obtain exclusive differential cross sections, a theoretical framework that accurately describes the hadron system is required. One such approach is based on the distorted-wave impulse approximation (DWIA) in the mean field approximation  [12, 13]. Mean-field-based models describe nucleons within an averaged nuclear potential generated by all other nucleons. This naturally leads to a shell structure, where each nucleon in the initial state is described by a bound wave function. Meanwhile, the distorted-wave treatment goes beyond the plane-wave description by also considering the final-state interactions of the outgoing nucleon: instead of emerging as a free particle, the outgoing nucleon propagates through the same mean field potential and undergoes distortions due to its interactions with the residual nucleus.
Despite their advantages, mean-field-based cross section calculations are computationally expensive compared to the factorized PWIA. The slow evaluation speed of these cross sections translates into slow sampling rates when traditional methods like accept-reject algorithms are used. To mitigate this, some current solutions [14, 15, 16, 17] involve precomputing the hadron tensor components into heavy multidimensional grids, known as hadron tables [8, 18, 19, 20]. This method allows to compute the cross section much faster, and make an accept-reject sampling possible. However, this approach has significant drawbacks: it lacks flexibility, requiring recalculation for any change in theoretical parameters, and seem infeasible for multi-nucleon knockout exclusive cross sections due to the exponential scaling of table size with dimensionality.
Recent advancements in Artificial Intelligence, particularly in Deep Learning, offer promising solutions to these challenges. In this work, we propose an alternative sampling method based on normalizing flows (NF) [21]. Normalizing flows transform a simple base distribution, such as a Gaussian, into a complex target distribution through a series of learnable diffeomorphisms. This enables efficient sampling and evaluation of the target distribution—in this case, the neutrino-nucleus cross section. In this work, we demonstrate the strong potential of NF to build a Monte Carlo event generator for neutrino-induced cross sections. For our proof-of-principle we use the Hartree-Fock mean-field exclusive 1-particle-1-hole ($1p1h$) cross section formalism developed by the Ghent group [22, 23, 13]. These cross sections exhibit all the features needed to demonstrate that our approach can pave the way for faster and more flexible event generation in neutrino physics.
@@ II Neutrino-nucleus exclusive CCQE cross section in the Mean Field Framework
The aim of this study is to demonstrate the ability to efficiently and accurately generate events for single nucleon knockout distributed according to the Hartree-Fock calculations developed by the Ghent group. While this model can predict both single and multiple nucleon knockout, providing a Monte Carlo event generator for all processes in their full complexity lies beyond the scope of this work. Future steps for handling the other processes are discussed in Section V. Here, we focus on $1p1h$ interactions due to a single-nucleon operator. The goal of this section is to introduce the relevant kinematic structure of the cross section.
@@
where $\omega=E_{\nu}-E_{\mu}$ is the energy transfer, $E_{N}$ the outgoing nucleon energy, and $\Omega$ the solid angles for the muon and proton. $E_{N}$ is the energy of the outgoing nucleon. The lepton tensor is independent of the nuclear model, and is calculated in electroweak theory. As the cross section is not differential in neutrino energy, event generators typically treat $E_{\nu}$ as an external input, modeling the conditional cross section for a fixed neutrino energy.
@@
We make an additional simplification to reduce the number of independent variables which is specific to the mean-field model. The hadron tensor $W^{\mu\nu}$ is a bilinear product of the hadron currents with an energy conserving delta function
@@
We thus generate samples of the four independent continuous variables $(\theta_{\mu},\omega,\theta_{N},\phi_{N})$ for each shell separately, with $E_{\nu}$ as an input. Among these variables, only the energy transfer $\omega$ has a more complex marginal distribution, particularly at low values where nuclear effects play a significant role. In contrast, the remaining variables are angles which are characterized by fixed support and exhibit relatively smooth distributions.
@@ III.1 Definition of normalizing flows
Normalizing Flows (NF) are a class of machine learning models used to transform simple probability distributions into complex ones through a sequence of invertible and differentiable transformations of the probability space. They are particularly useful for efficiently sampling from high-dimensional and non-Gaussian distributions, making them highly suited for generating events from complex interaction models, such as neutrino-nucleus interactions.
The core idea behind NF is to start with a simple base distribution, typically a multidimensional Gaussian or uniform distribution, and apply a series of transformations that map this distribution into the more complex target distribution of interest. Mathematically, a typical NF transformation consists of a sequence of $K$ diffeomorphisms of the probability space, each parameterized by a neural network, that progressively transform the base distribution $p_{Z}(z)$ into the target distribution $p_{X}(x)$. If $z\sim p_{Z}(z)$ is a sample from the base distribution, the final output $x=f_{K}\circ f_{K-1}\circ\cdots\circ f_{1}(z)$ represents a sample from the target distribution $p_{X}(x)$.
@@ III.5 Exhaustive training of the flow-based event generator
can lead to mean-seeking behavior, where the learned distribution $q_{\theta}$ covers the support of the true distribution $p$ but may overestimate the distribution tails. Conversely, minimizing the reverse KL-D, expressed as
@@
can result in mode-seeking behavior, where $q_{\theta}$ focuses on some modes of $p$ but may ignore other regions.
To balance these tendencies, we use the symmetric KL-D, also known as Jensen-Shannon metric, which is the average of these two:
@@ IV.1 Detail of implementation and training
We trained a flow-based Monte Carlo event generator for the CCQE $1p1h$ interaction between a neutrino and a carbon nucleus (${}^{12}\text{C}$). The carbon nucleus consists of two shells, $1s_{\nicefrac{{1}}{{2}}}$ and $1p_{\nicefrac{{3}}{{2}}}$, with respective separation energies of $36.20$ MeV and $18.72$ MeV. To enhance the flexibility of our event generation, we chose to model the interactions for each shell separately. This approach serves theoretical purposes, enabling the study of interactions with individual shells. A similar work can be applied to any nuclei, which would involve modeling more or less shells.
@@
We used a relatively low learning rate of $0.0001$ with a Cosine annealing scheduler with a minimum learning rate of $5\times 10^{-7}$. We trained each model for $100000$ epochs. Each epoch computes the Symmetric KL-D on a batch of $2^{14}$ samples with associated true cross section. $10$ million samples were used to train each model in total. These samples and their cross sections were pre-generated prior to training—a potentially time-intensive step depending on the cross section computation speed, but required only once. The $\theta_{\mu}$ and $\theta_{N}$ distributions are unfolded to $[-\pi,\pi]$ to ensure $2\pi$-periodicity. The training time for each model (on an NVIDIA RTX-4090 GPU) are given in Table 1.
@@ IV.2 Comparison of the true cross section and its surrogate projecting in 1D and 2D spaces
Figure 2 illustrates the comparison between the true cross section and its surrogate modeled using normalizing flows for a fixed neutrino energy of $E_{\nu}=200$ MeV, with the initial bound nucleon in the $1p_{\frac{3}{2}}$ shell. This specific combination was selected due to the pronounced influence of nuclear effects on its cross section. At low energy transfer, the cross section exhibits a pronounced dependence on shell structure effects, a core addition of theoretical nuclear shell models such as the Hartree-Fock mean-field approach. The angular distribution of the outgoing nucleon relative to the transferred momentum, $\theta_{N}$, is strongly influenced by the properties of the initial bound state. Nevertheless, the NF seem to model accurately this distribution from the shapes of the four marginal distributions to the correlations between two kinematic variables.
@@ IV.3 Comparison of the true cross section and its surrogate using density estimation
where $\alpha$ refers to the bound nucleon shell.
@@ IV.4 Comparison between the true cross section and its surrogate through their respective datasets
In neutrino experiments such as T2K, we cannot achieve exact estimates of the kinematics we measure or infer. As a result, evaluating a model based on a direct point-by-point comparison of its distribution with the true one, as discussed in Section IV.3, may be overly conservative. Instead, the goal in this work is to construct an event generator capable of populating the kinematic phase space in a manner consistent with a non-biased accept-reject method. This must be achieved with fixed precision while accounting for the full dimensionality of the phase space.
@@
To provide a measure of how the surrogate’s MNLL values differ from those of the accept-reject method in average, we compute the averaged Z-Score (Z-S) across all energies for a given shell:
$\begin{split}\text{Z-Score}=\frac{1}{N_{\text{e}}}\sum_{i=1}^{N_{\text{e}}}\left(\frac{\overline{\text{MNLL}(H^{\text{NF}}_{E_{\nu}^{(i)}})}-\overline{\text{MNLL}(H^{\text{AR}}_{E_{\nu}^{(i)}})}}{\sigma^{\text{AR}}_{E_{\nu}^{(i)}}}\right).\end{split}$
[…]
@@ IV.5 Overall performance
Z-Score ($1s_{\nicefrac{{1}}{{2}}}$) 1.17 0.92 0.66
Z-Score $(1p_{\nicefrac{{3}}{{2}}}$) 1.66 1.32 0.82
@@
The results show that models with larger sizes provide higher accuracy, as indicated by a steady rise in RESS values from 98.48% (S) to 98.87% (L). However, this improvement is relatively small, suggesting that the smallest model might be sufficient when sampling speed or memory constraints outweigh minor accuracy gains. This trade-off between sampling speed and accuracy depends on how the surrogate cross section is utilized.
@@
    Sample directly from the surrogate without any further resampling. This option is by far the fastest for cross sections that are computationally heavy. However, it is important to mention that no generative model can replace a true cross section without introducing bias. Therefore, for a strictly unbiased analysis, one should always rely on the importance weight.
@@
Moreover, our normalizing flows approach to build a proposal distribution already achieves greater accuracy than the previous normalizing flows applied to neutrino-nucleus cross section modelling: Pina-Otey et al. [46] reported a RESS of 91.40%, which is significantly lower than even our smallest model (S). The concentration of weights around $1$ as shown in Figure 3 highlights the efficiency of the surrogate cross section as a proposal distribution for an accept-reject algorithm. The low spread of weights ensures that most samples contribute effectively to the estimation of observables, minimizing the variance introduced by the importance weights. If the weight threshold is set at $1.90$, corresponding to the $99.99^{\text{th}}$ percentile of weights then only one in ten thousand samples surpasses this limit. The expected acceptance rate for the surrogate cross section with this weight cap, when used as a proposal distribution, is approximately $\frac{1}{1.90}\approx 53\%$.
Although the surrogate can serve as an efficient proposal distribution, Section IV.4 shows that a dataset of events sampled directly from the normalizing flow surrogate is practically equivalent to an unbiased accept-reject dataset at our target precision. Consequently, sampling directly from the surrogate cross-section, therefore using NF as an emulator, can be used depending on the accuracy we require or at first approximation. This is particularly advantageous for complex, theory-driven cross sections requiring lengthy evaluations, since normalizing flows can reduce total sampling time from several days on a CPU (via accept-reject) to a few minutes on a GPU or CPU.

### v3

@@ I Introduction
To obtain exclusive differential cross sections, a theoretical framework that accurately describes the hadron system is required. One such approach is based on the distorted-wave impulse approximation (DWIA) in the mean field approximation  [12, 13]. Mean-field-based models describe nucleons within an averaged nuclear potential generated by all other nucleons. This naturally leads to a shell structure, where each nucleon in the initial state is described by a bound wave function. Meanwhile, the distorted-wave treatment goes beyond the plane-wave description by also considering the final-state interactions of the outgoing nucleon: instead of emerging as a free particle, the outgoing nucleon propagates through the same mean field potential and undergoes distortions due to its interactions with the residual nucleus. This approach has recently gained renewed attention in the context of event generator development, with the first implementation of a relativistic DWIA-based CCQE model in NEUT [14].
Despite their advantages, mean-field-based cross section calculations are computationally expensive compared to the factorized PWIA. The slow evaluation speed of these cross sections translates into slow sampling rates when traditional methods like accept-reject algorithms are used. To mitigate this, some current solutions [15, 16, 17, 18] involve precomputing the hadron tensor components into heavy multidimensional grids, known as hadron tensor tables [8, 19, 20, 21]. This method allows to compute the cross section much faster, and make an accept-reject sampling possible. However, this approach has significant drawbacks: it lacks flexibility, requiring recalculation for intricate change in theoretical parameters, and seem infeasible for multi-nucleon knockout exclusive cross sections due to the exponential scaling of table size with dimensionality.
Recent advancements in Artificial Intelligence, particularly in Deep Learning, offer promising solutions to these challenges [22, 23, 24, 25]. In this work, we propose an alternative sampling method based on normalizing flows (NF) [26]. Normalizing flows transform a simple base distribution, such as a Gaussian, into a complex target distribution through a series of learnable diffeomorphisms. This enables efficient sampling and evaluation of the target distribution—in this case, the neutrino-nucleus cross section. In this work, we demonstrate the capacity of NF to build a Monte Carlo event generator for neutrino-induced cross sections. For our proof-of-principle we use the Hartree-Fock mean-field exclusive 1-particle-1-hole ($1p1h$) cross section formalism developed by the Ghent group [27, 28, 13]. These cross sections exhibit all the features needed to demonstrate that our approach can pave the way for faster and more flexible event generation in neutrino physics.
@@ II Neutrino-nucleus exclusive CCQE cross section in the Mean Field Framework
The aim of this study is to demonstrate the ability to generate events with precision and speed for single nucleon knockout distributed according to the Hartree-Fock calculations developed by the Ghent group. While this model can predict both single and multiple nucleon knockout, providing a Monte Carlo event generator for all processes in their full complexity lies beyond the scope of this work. Future steps for handling the other processes are discussed in Section V. Here, we focus on $1p1h$ interactions due to a single-nucleon operator. The goal of this section is to introduce the relevant kinematic structure of the cross section.
@@
where $\omega=E_{\nu}-E_{\mu}$ is the energy transfer, $E_{N}$ the outgoing nucleon energy, and $\Omega$ the solid angles for the muon and proton. The lepton tensor is independent of the nuclear model, and is calculated in electroweak theory. As the cross section is not differential in neutrino energy, event generators typically treat $E_{\nu}$ as an external input, modeling the conditional cross section for a fixed neutrino energy.
@@
We make an additional simplification (without incurring any loss of information) to reduce the number of independent variables which is specific to the mean-field model. The hadron tensor $W^{\mu\nu}$ is a bilinear product of the hadron currents with an energy conserving delta function :
@@
We thus generate samples of the four independent continuous variables $(\theta_{\mu},\omega,\theta_{N},\phi_{N})$ for each shell separately, with $E_{\nu}$ as an input. Among these variables, only the energy transfer $\omega$ has a more complex marginal distribution, particularly at low values where nuclear effects play a significant role. This is due to the Hartree-Fock method preserving the shell structure of the nucleus which in return impacts the cross section. In contrast, the remaining variables are angular and exhibit relatively smooth distributions over well-defined, bounded ranges.
@@ III.1 Definition of normalizing flows
Normalizing Flows (NF) are a class of machine learning models used to transform simple probability distributions into complex ones through a sequence of invertible and differentiable transformations of the probability space. They are particularly useful for efficiently sampling from high-dimensional and non-Gaussian distributions, making them highly suited for generating events from complex interaction models, such as neutrino-nucleus interactions. This contrasts with the accept-reject loop used in most neutrino event generators, which discards a large fraction of generated samples.
The core idea behind NF is to start with a simple base distribution, typically a multidimensional Gaussian or uniform distribution, and apply a series of transformations that map this distribution into the more complex target distribution of interest. Mathematically, a typical NF transformation consists of a sequence of $K$ diffeomorphisms of the probability space, each parameterized by a neural network, that progressively transform the base distribution $p_{Z}(z)$ into the target distribution $p_{X}(x)$. If $z$ is a sample from the base distribution $p_{Z}(z)$, the final output $x=f_{K}\circ f_{K-1}\circ\cdots\circ f_{1}(z)$ represents a sample from the target distribution $p_{X}(x)$.
@@ III.5 Exhaustive training of the flow-based event generator
can lead to mean-seeking behavior, where the learned distribution $q_{\theta}$ covers the regions where the true distribution $p$ is non-zero but may overestimate the distribution tails. Conversely, minimizing the reverse KL-D, expressed as
@@
can result in mode-seeking behavior, where $q_{\theta}$ focuses on some peaks in the probability but may ignore other regions or even other peaks.
To balance these tendencies, we use the symmetric KL-D, which is the average of these two:
@@ IV.1 Detail of implementation and training
We trained a flow-based Monte Carlo event generator for the CCQE $1p1h$ interaction between a neutrino and a carbon nucleus (${}^{12}\text{C}$). The carbon nucleus consists of two shells, $1s_{\nicefrac{{1}}{{2}}}$ and $1p_{\nicefrac{{3}}{{2}}}$, with respective separation energies of $36.20$ MeV and $18.72$ MeV. To enhance the flexibility of our event generation, we chose to model the interactions for each shell separately. This approach serves theoretical purposes, enabling the study of interactions with individual shells. A similar work can be applied to any nuclei, which would involve modeling more or fewer shells.
@@
We used a relatively low learning rate of $0.0001$ with a Cosine annealing scheduler with a minimum learning rate of $5\times 10^{-7}$. We trained each model for $100000$ epochs. Each epoch computes the Symmetric KL-D on a batch of $2^{14}$ samples with associated true cross section. A total of $10$ million samples were used to train each model, with each sample appearing multiple times throughout the training. These samples and their cross sections were pre-generated prior to training—a potentially time-intensive step depending on the cross section computation speed, but required only once. The $\theta_{\mu}$ and $\theta_{N}$ distributions are unfolded to $[-\pi,\pi]$ to ensure $2\pi$-periodicity. The training time for each model (on an NVIDIA RTX-4090 GPU) are given in Table 1.
@@ IV.2 Comparison of the true cross section and its surrogate projecting in 1D and 2D spaces
Figure 2 illustrates the comparison between the true cross section and its surrogate modeled using normalizing flows for a fixed neutrino energy of $E_{\nu}=200$ MeV, with the initial bound nucleon in the $1p_{\frac{3}{2}}$ shell. This particular configuration was chosen because nuclear effects have a significant impact on its cross section. At low energy transfer, the cross section is especially sensitive to shell structure, a key feature captured by the Hartree-Fock mean-field approach. The angular distribution of the outgoing nucleon relative to the transferred momentum, $\theta_{N}$, is strongly influenced by the properties of the initial bound state. Nevertheless, the NF seem to model accurately this distribution from the shapes of the four marginal distributions to the correlations between two kinematic variables.
@@ IV.3.1 Point-by-point density comparison
where $\alpha$ refers to the bound nucleon shell. This is the error one expects in a sample if the NF surrogate was used in its simplest form, and also the weight one would want to apply to an event generator sample to correct the imperfect surrogate to be closer to the true distribution.
@@
IV.3.2 Density comparison with finite resolution
Given that the RESS value obtained is slightly below $1$, it is necessary to assess the potential impact of the observed weight variation. A central question is whether the largest deviations are concentrated in specific regions of the phase space, which could introduce systematic biases in the generated sample exceeding current experimental uncertainties. Alternatively, these variations may stem from localized fluctuations which, when integrated over sufficiently small kinematic bins, effectively average out and thus remain negligible for experimental observations.
To clarify this point, the true and surrogate cross sections are both evaluated on a grid for a monoenergetic neutrino beam at $650$ MeV, summing both shell contributions. The grid spans different muon scattering angles with a step of $10^{\circ}$ and outgoing proton angles, specifically $\theta_{N}\in{1^{\circ},5^{\circ},10^{\circ},20^{\circ},30^{\circ},40^{\circ},50^{\circ},60^{\circ}}$ with a fixed nucleon azimuthal angle of $180^{\circ}$. The energy transfer $\omega$ is sampled with a step of $2.5$ MeV across its defined range. Importantly, the cross section values are averaged within localized four-dimensional kinematic bins centered at each point $(\omega,\theta_{\mu},\theta_{N},\phi_{N})$. The resolution of each bin is $1$ MeV in energy transfer and $1^{\circ}$ in angular coordinates, which is significantly finer than what current experiments can resolve. Every bin includes $500$ uniformly distributed samples, ensuring negligible statistical fluctuations in the averaged bin value.
Figure 4 presents the histogram of weights after averaging, of the given sampled points weighted by the true cross section values. The dashed lines represent the $68\%$, $95\%$, and $99\%$ confidence intervals. Appendix A further illustrates the impact of these cross section deviations on the distribution of proton kinetic energy $T_{N}$ at fixed kinematics, both with and without local averaging.
To quantify the residual discrepancies between the surrogate and true cross sections after local binning, we introduce the cross section-weighted root mean squared error (RMSp), defined as:
$\text{RMS}_{p}=\sqrt{\frac{\sum_{\mathbf{x}_{i}}p(\mathbf{x}_{i})\,\Delta^{2}(\mathbf{x}_{i})}{\sum_{\mathbf{x}_{i}}p(\mathbf{x}_{i})\,}},$
where:
  • $\mathbf{x}_{i}$ denotes a point in the kinematic space,
  • $\Delta(\mathbf{x}_{i})$ is the relative difference between the surrogate and true cross section at point $\mathbf{x}_{i}$ after averaging,
  • $p(\mathbf{x}_{i})$ is the true cross section at $\mathbf{x}_{i}$ after averaging.
We obtain an $\text{RMS}_{p}$ of $1.4\%$. This demonstrates that the local averaging procedure drastically reduces the modeling error compared to the initial spread of weights of $10.7\%$. The bias introduced by Normalizing Flows is therefore well below our experimental uncertainties. Furthermore, in realistic experimental conditions, where measurements involve a convolution over a broad neutrino energy spectrum, such residual differences would be further smeared out, rendering them more negligible in practice.
Figure 4: Histogram of the ratio between true and surrogate cross sections weighted by the true cross section evaluated at a neutrino energy of $650$ MeV across a fine kinematic grid and after local averaging. The dashed lines represent the confidence intervals: $68\%$ CI = ($-0.75\%$, $+0.69\%$), $95\%$ CI = ($-1.88\%$, $+1.92\%$), and $99\%$ CI = ($-4.00\%$, $+3.57\%$).
@@ IV.4 Comparison between the true cross section and its surrogate through their respective datasets
In neutrino experiments such as T2K, we cannot achieve exact estimates of the kinematics we measure or infer. As a result, evaluating a model based on a direct point-by-point comparison of its distribution with the true one, as discussed in Section IV.3.2, may be overly conservative. Instead, the goal in this work is to construct an event generator capable of populating the kinematic phase space in a manner consistent with a non-biased accept-reject method. This must be achieved with fixed precision while accounting for the full dimensionality of the phase space.
@@
To provide a measure of how the surrogate’s MNLL values differ from those of the accept-reject method in average, we compute the RMS of the Z-Score ($\text{RMS}_{z}$) across all energies for a given shell:
$\begin{split}\text{RMS}_{z}=\sqrt{\frac{1}{N_{\text{e}}}\sum_{i=1}^{N_{\text{e}}}\left(\frac{\overline{\text{MNLL}(H^{\text{NF}}_{E_{\nu}^{(i)}})}-\overline{\text{MNLL}(H^{\text{AR}}_{E_{\nu}^{(i)}})}}{\sigma^{\text{AR}}_{E_{\nu}^{(i)}}}\right)^{2}}.\end{split}$
where $\overline{\text{MNLL}(H^{\text{NF}}_{E_{\nu}^{(i)}})}$ is the mean MNLL for normalizing flow datasets at energy $E_{\nu}^{(i)}$, $\overline{\text{MNLL}(H^{\text{AR}}_{E_{\nu}^{(i)}})}$ is the mean MNLL for accept-reject datasets at the same energy, and $\sigma^{\text{AR}}_{E_{\nu}^{(i)}}$ is the standard deviation of the accept-reject MNLL distribution at energy $E_{\nu}^{(i)}$. The $\text{RMS}_{z}$ for the three model sizes are given in Table 1. The MNLL distributions for the surrogate and accept-reject histograms are statistically compatible within their $\text{RMS}_{z}=1$ range for the $L$ model and within $\text{RMS}_{z}=2$ for the $M$ and $S$ models. In the case of the model size L, we find an average $\text{RMS}_{z}$ of $0.75\,\sigma$ for the $1s_{\nicefrac{{1}}{{2}}}$ shell and $0.88\,\sigma$ for the $1p_{\nicefrac{{3}}{{2}}}$ shell. In other words, two datasets of $100,000$ samples respectively generated by accept-reject and by the normalizing flow surrogate are practically equivalent given the binning that we chose. Overall, these results demonstrate that the normalizing flow surrogate reproduces the cross section with high fidelity in all tested combinations of shell and energy and in the full kinematic space, while significantly reducing computational overhead compared to the accept-reject method.
@@ IV.5 Overall performance
$\text{RMS}_{z}$ ($1s_{\nicefrac{{1}}{{2}}}$) 1.36 1.01 0.75
$\text{RMS}_{z}$ $(1p_{\nicefrac{{3}}{{2}}}$) 1.86 1.52 0.88
@@
The results show that models with larger sizes provide higher accuracy, as indicated by a steady rise in RESS values from 98.48% (S) to 98.87% (L). However, this improvement is relatively small in terms of unweighting efficiency in the case of important sampling, suggesting that the smallest model might be sufficient when sampling speed or memory constraints outweigh minor accuracy gains in the proposal distribution. This trade-off between sampling speed and accuracy depends on how the surrogate cross section is utilized.
@@
    Sample directly from the surrogate without any further resampling. This option is by far the fastest for cross sections that are computationally heavy. However, it is important to mention that no generative model can replace a true cross section without introducing bias. Therefore, for a strictly unbiased analysis, one should always rely on the importance weight. As mentioned in Section IV.3, this is particularly crucial for studies focused on regions of very low cross section (e.g., low proton kinetic energy or backward-going protons with respect to the momentum transfer) or when very fine resolution is needed, such as angular resolutions better than one degree or energy resolutions finer than $1$ MeV.
@@
Moreover, our normalizing flows approach to build a proposal distribution already achieves greater accuracy than the previous normalizing flows applied to neutrino-nucleus cross section modeling: Pina-Otey et al. [50] reported a RESS of $91.40\%$, which is significantly lower than even our smallest model (S). The concentration of weights around $1$ as shown in Figure 3 highlights the efficiency of the surrogate cross section as a proposal distribution for an accept-reject algorithm. The low spread of weights ensures that most samples contribute effectively to the estimation of observables, minimizing the variance introduced by the importance weights. If the weight threshold is set at $1.90$, corresponding to the $99.99^{\text{th}}$ percentile of weights then only one in ten thousand samples surpasses this limit. The expected acceptance rate for the surrogate cross section with this weight cap, when used as a proposal distribution, is approximately $\frac{1}{1.90}\approx 53\%$.
Although the surrogate can serve as an efficient proposal distribution, Section IV.4 shows that a dataset of events sampled directly from the normalizing flow surrogate is practically equivalent to an unbiased accept-reject dataset at our target precision. Section IV.3 and Appendix A further show that, once experimental conditions are taken into account, particularly a reasonable resolution in observables, the resulting uncertainties are already at the percent level for a monoenergetic neutrino beam. This residual error would be further reduced after convolution with a realistic neutrino flux, due to the additional smearing over neutrino energies. Importantly, the remaining modeling uncertainty is already well below the typical systematic uncertainties associated with experimental measurements, making the Normalizing Flows surrogate sufficiently accurate for practical analyses. This is particularly advantageous for complex, theory-driven cross sections requiring lengthy evaluations, since normalizing flows can reduce total sampling time from several days on a CPU (via accept-reject) to a few minutes on a GPU or CPU.